OpenStudy (anonymous):

Help again please! :D Evaluate the definite integral... from 0 to pi 2sin(x) dx

OpenStudy (anonymous):

you know the anti derivative of sine?

OpenStudy (anonymous):

oh...is it not asking for the antiderivative of the antiderivative??? if i am saying that right lol

OpenStudy (anonymous):

antiderivative of sine would be -cos right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your answer is therefore \[-2\cos(\pi)+2\cos(0)\]

OpenStudy (anonymous):

i am just confused to why the answer is just the antiderivative

OpenStudy (anonymous):

and wouldn't you want to subtract the values from each other not add them? or is that from the derivative of the antiderivative?